Solve for x.
| 2x - 5 | + 6 = 12
need answer please :)
: Solve for x.
| 2x - 5 | + 6 = 12
need answer please :)

Answer by stanbon(19007)

| 2x - 5 | + 6 = 12
|2x-5| = 6
2x-5 = 6 or 2x-5 = -6
2x = 11 or 2x = -1
x = 11/2 or x = -1/2
========================
Cheers,
Stan H.